OpenStudy (anonymous):

Help Please : ( Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1).

OpenStudy (anonymous):

you need the equations parallel and perpendicular to y = -3x - 1, y=mx+ b --> slope intercept form, m=slope, b=y-int \[(y-y _{1})=m(x-x _{1})\]--> equation of the line where m is the slope, (x1, y1) is any point on the line.. first you need to find the slopes of the lines: slope of parallel lines is the same as the slope of the original equation slope of perpendicular line is the negative reciprocal of the slope of the original equation then using the the given point (3, 1) plug it to the equation of the line with the corresponding slope and simplify.

OpenStudy (anonymous):

parallel: y = -3x + 10 perpendicular: y = one over threex @Data_LG2 ?

OpenStudy (anonymous):

give me a sec. i'll try to do it

OpenStudy (anonymous):

lol, i made a mistake, wait

OpenStudy (anonymous):

 YEAH YOU ARE RIGHT @MathHater1994